# Enhancing Financial Time-Series Forecasting with Retrieval-Augmented Large Language Models

### v2

@@
Abstract.
Stock movement prediction, a critical task in financial time-series forecasting, relies on identifying and retrieving key influencing factors from vast and complex datasets. However, traditional text-trained or numeric similarity-based retrieval methods often struggle to handle the intricacies of financial data. To address this, we propose the first retrieval-augmented generation (RAG) framework specifically designed for financial time-series forecasting. Our framework incorporates three key innovations: a fine-tuned 1B large language model (StockLLM) as its backbone, a novel candidate selection method enhanced by LLM feedback, and a training objective that maximizes the similarity between queries and historically significant sequences. These advancements enable our retriever, FinSeer, to uncover meaningful patterns while effectively minimizing noise in complex financial datasets. To support robust evaluation, we also construct new datasets that integrate financial indicators and historical stock prices. Experimental results demonstrate that our RAG framework outperforms both the baseline StockLLM and random retrieval methods, showcasing its effectiveness. FinSeer, as the retriever, achieves an 8% higher accuracy on the BIGDATA22 benchmark and retrieves more impactful sequences compared to existing retrieval methods. This work highlights the importance of tailored retrieval models in financial forecasting and provides a novel, scalable framework for future research in the field.
Retrieval-augmented generation, Financial Time-series forecasting, Stock Movement Prediction, Large Language Model
copyright: noneccs: Information systems Top-k retrieval in databasesccs: Information systems Similarity measuresccs: Information systems Novelty in information retrieval
1. Introduction
Financial time-series forecasting plays a pivotal role in ensuring market stability and efficiency, directly impacting critical areas such as investment strategies, risk management, and the formulation of economic policies (Fama and French, 2000). One of the most prominent tasks in this domain is stock movement prediction (Xie et al., 2023a, 2024b, 2024a), which focuses on forecasting the direction of price changes, i.e. whether a stock will rise or fall. Accurate predictions of financial metrics, including stock movements, interest rates, and economic indicators, are essential for a wide range of stakeholders, from individual investors to large financial institutions. However, the inherent complexity and volatility of financial markets pose significant challenges to forecasting, necessitating the use of advanced methodologies to effectively analyze vast and often noisy datasets.
Traditional stock movement prediction methods (Feng et al., 2018; Qin et al., 2017) relied on analyzing sequences of stock prices and financial indicators [^1] Financial indicators (Jansen, 2020) are professional market analysis metrics derived from price data, where their values or crossover points signal fluctuations in the stock market, indicating rising or falling trends. However, relying on one or two indicators can introduce bias, and if an indicator only provides meaningful information at crossover points, it may fail to offer valuable insights during periods without crossings. to identify patterns, but they often failed to capture the semantic connections between these sequences, a task at which large language models (LLMs) excel (Huang and Chang, 2022). However, LLM-based methods (Xie et al., 2023b, 2024a, a) have primarily shifted focus to leveraging textual data such as news and tweets, often using only the past several days’ closing prices as reference while overlooking the rich insights embedded in historical time-series data (Bustos and Pomares-Quimbaya, 2020). This highlights the need to effectively integrate financial time-series data with LLMs. Directly working with such data presents a challenge due to its vast scale, encompassing both the breadth of influencing variables and the depth of historical trends. To address this, a retrieval-based approach is necessary to efficiently sift through extensive time-series data, extracting meaningful information to enhance LLMs’ ability to deliver more accurate and robust stock movement predictions.
Although retrieval-augmented generation (RAG) methods have been applied for various tasks (Joshi et al., 2024; He et al., 2024; Cui et al., 2024), it is challenging to directly apply them to financial time-series retrieval. First, most embedding-based retrieval methods (retrievers) (Su et al., 2023; Li and Li, 2023; Xiao et al., 2023; Zhang et al., 2023; Wang et al., 2023) are trained on textual data and struggle to process numeric time-series data, which lacks explicit semantic information. Second, distance-based retrieval methods like Dynamic Time Warping (DTW) (Yang et al., 2024) focus on numeric similarities and fail to capture deeper semantic relationships or contextual nuances essential for stock movement prediction. These limitations underscore the need for a more tailored retrieval framework designed for financial time-series data.
[…]
The framework of FinSeer, illustrating its components and workflow.
Figure 1. Overview of our time-series RAG framework.
To address these challenges, we introduce a retrieval-augmented framework for stock movement prediction, featuring the first dedicated retriever for financial time-series forecasting, Financial Time-Series Retriever (FinSeer). Unlike previous RAG methods, which rely on pre-trained encoders or distance-based metrics, our framework is the first to train a dedicated retriever tailored for continuous and complex temporal sequences. The whole retrieval framework is shown in Figure 1.
To address the limitations of existing datasets, which often contain only price data, and to better align with professional financial analysis, we construct new datasets by integrating 20 groups of financial indicators. These indicators provide critical insights into market behavior that stock prices alone cannot capture [^2] For instance, the overbought area in the KDJ indicator group (Wu and Diao, 2015), where $K>80$, $D>70$, and $J>90$, signals a potential falling trend.. We begin by selecting high-trade-volume U.S. stocks, and the indicators are chosen through domain expert consultation and mutual information analysis. These indicators are then serialized and divided into train, valid and test sets. By training the retriever on datasets enriched with diverse financial indicators, our retrieval framework incorporates financial domain expertise, enabling it to interpret the implications of trending signals that uncover deeper patterns beyond surface-level trend similarities. The testing set, designed to mimic real-world economic analysis, ensures a robust evaluation of model performance in practical financial scenarios.
To train our retriever to identify semantically related sequences and align the retriever’s priorities with those of the LLM, we introduce a novel retrieval mechanism with three key innovations: a new backbone LLM, a novel method for selecting positive and negative candidates, and a new training objective. First, to activate the LLM’s inherent knowledge and ensure instruction-following capabilities, we fine-tune a 1B parameter LLM (LLaMA3.2-1B Instruct) (Grattafiori et al., 2024), creating StockLLM as the backbone model of our retriever. By using a smaller LLM, we establish a more challenging experimental setup, ensuring performance improvements are attributable to FinSeer rather than the LLM’s capacity. Next, StockLLM identifies time-series segments that enhance the generative process, feedbacks from StockLLM are used to select the most beneficial sequences that could lead the model to make the correct decision as positive candidates and the least beneficial as negative candidates. This could bridge the gap between the retriever and the LLM and ensure that retrieved data aligns with the LLM’s forecasting priorities. Finally, the training objective, inspired by Zhang et al. (2023), ensures the retriever distinguishing historically significant sequences (positive candidates) from noisy sequences (negative candidates) by maximizing the similarity between the query and candidate sequences. This prioritizes meaningful patterns while minimizing irrelevant data, outperforming traditional methods that struggle to separate signal from noise in complex financial time series.
Based on FinSeer, the inference process of our RAG framework consists of three steps: (1) using FinSeer to retrieve relevant sequences from the candidate pool, (2) incorporating these sequences into StockLLM’s input context, and (3) combining this enriched context with the original query to perform forecasting. This approach effectively integrates salient historical patterns with current data, improving prediction accuracy.
To thoroughly evaluate the performance of our RAG framework and the retriever, we rely solely on temporal information, excluding additional textual data like news. For the RAG framework evaluation, StockLLM serves as the backbone model, integrated with different retrieval methods for comparison. Experimental results demonstrate that using our RAG framework outperforms bare StockLLM, indicating that the RAG framework enhances LLM performance by incorporating relevant time-series data as financial context. However, RAG with retrieval models trained on text data does not always improve performance compared to bare StockLLM. In some cases, performance even declines. This suggests that these retrieval models cannot effectively identify beneficial time-series data that could enhance the model’s predictions, which highlights a fundamental gap between text-based retrieval and time-series data. Furthermore, RAG framework with FinSeer consistently surpasses all other retrieval models, such as Instructor (Su et al., 2023), BGE (BAAI General Embedding) (Xiao et al., 2023), LLM-Embedder (Zhang et al., 2023), UAE (Li and Li, 2023), and E5-mistral-7b-instruct (Wang et al., 2023), demonstrating the effectiveness of our trained FinSeer in identifying relevant time-series financial data to enhance stock movement prediction. Further analysis of retrieval results reveals that FinSeer uniquely identifies relevant financial indicators that enhance LLM performance, while other retrievers predominantly select sequences based on superficial similarities, such as close price and adjusted close price. This behavior underscores a critical limitation of text-trained models: they focus on surface-level patterns and fail to discern truly valuable time-series information. In contrast, FinSeer effectively bridges this gap, demonstrating its ability to uncover underlying relevance in financial data.
Our contributions are summarized as follows:
  1. (1)
    We propose the first RAG framework for financial time-series forecasting, in which the retrieval mechanism features three key innovations: a fine-tuned 1B parameter LLM (StockLLM) as the backbone, a novel candidate selection method leveraging LLM feedback to identify beneficial sequences, and a training objective that maximizes similarity between queries and historically significant sequences, enabling the retriever to uncover meaningful patterns while minimizing noise in complex financial data.
  2. (2)
    We construct new training and testing datasets to address the limitations of existing ones containing only price data, and better align with real-world economic analysis. Besides price data, these datasets integrate financial indicators that signal upward or downward trends, carefully selected through domain expert consultation and mutual information analysis. By using new training datasets, we train the retriever to uncover deeper patterns beyond surface-level trends during training. While the testing set, designed to mimic real-world scenarios, ensures robust evaluation of model performance in practical financial applications.
  3. (3)
    We propose a new retriever, FinSeer, specifically designed for time-series forecasting, trained on newly constructed datasets that integrate financial indicators alongside historical stock prices. By leveraging LLM feedback and prioritizing historically significant sequences, FinSeer uncovers deeper patterns beyond surface-level trends, setting it apart from traditional methods.
  4. (4)
    Experimental results demonstrate that our RAG framework outperforms bare StockLLM and random retrieval, highlighting its effectiveness, while our retriever, FinSeer, surpasses existing RAG methods in stock movement prediction across three datasets, achieving an 8% higher accuracy on BIGDATA22 compared to a general-purpose LLM-feedback-based retriever. Furthermore, FinSeer retrieves more dynamic and impactful sequences, showcasing its ability to identify and leverage the most relevant patterns for forecasting.
2. Problem Definition and Goals
@@
Retrieval-augmented financial time-series forecasting (Sezer et al., 2020) involves predicting future values or trends ($G$) based on a given query sequence ($q$) and a set of retrieved historical sequences ($c$). These sequences are collected over time at regular intervals. In the retrieval process, the goal of the retrieval model ($R$) is to efficiently identify and extract the most useful historical sequences from a large pool of candidates. By providing relevant context, the retrieval model enhances the forecasting model’s ability to make accurate and reliable predictions.
Stock Movement Prediction Task. In the specific task of stock movement prediction (Xu and Cohen, 2018), the problem is framed as a binary classification task: predicting whether a stock’s price will rise or fall on the next trading day. Given a query sequence $q$, which represents the stock’s price over the previous $t$ days, the model retrieves relevant sequences as context and predicts the stock’s movement $M_{q,d}$ for the next trading day $d$. Table 1 defines the major symbols used in this paper.
Table 1. The definition of symbols.
General Time-series Symbol Stock Movement Prediction Symbol Definitions
$q$ $q=\{q_{d-t},...,q_{d-1}\}$ The query time-series data. In stock movement prediction, $q$ refers to the query stock price sequence of length $t$, containing stock price data from trading day $d-t$ to trading day $d-1$.
$G(q)$ $G(q,d)\in\{\textit{rise},\textit{fall}\}$ The final output $G$ given the query $q$. In stock movement prediction, $G(q,d)$ shows the generation $G$ of the query stock $q$ on the query trading day $d$, belonging to rise or fall.
$P(c)=LLM(O|q,c)$ $P(c)=LLM(M_{d}|q,c)$ The probability $P$ of the LLM to generate an accurate output $O$ given the query sequence $q$ and a candidate sequence $c$. In stock movement prediction, $P(c)$ refers to generating the accurate movement $M$ on the query trading day $d$.
=$\mathbb{C_{P}}=\{c_{i}\mid i=1,\ldots,k\}$ The set of top-k retrieved sequences as positive examples, where $P(c_{i})\geq P(c_{i+1})$.
=$\mathbb{C_{N}}=\{c_{i}\mid i=k+1,\ldots,n\}$ The set of negative retrieved sequences, where $P(c_{i})\geq P(c_{i+1})$.
$w_{i}$ The soft weight of the $i^{th}$ retrieved sequences, where $w_{i}=P(c_{i}),i=1,...,k$.
$R$ The retrieve model.
[…]
Following Yoo et al. (2021) and  Soun et al. (2022), we classify the movement as rise if return exceeds 0.55, fall if it is below -0.5. In line with previous experimental settings (Xie et al., 2023b, 2024a), we do not evaluate freeze cases. However, we include sequences with $R_{t}\in[-0.5,0.55]$ as freeze candidates in the candidate pool, ensuring a diverse and comprehensive set of historical data for context.
3. The RAG Framework
To identify valuable sequences from vast amounts of stock data and conduct stock movement prediction, we propose a novel RAG framework with new datasets and a novel retrieval mechanism with a new backbone LLM. First, we construct three datasets that integrate financial indicators alongside historical stock prices, addressing the limitations of existing datasets and better aligning with real-world economic analysis. Second, we introduce StockLLM, a fine-tuned 1B parameter LLM, as the backbone to ensure instruction-following capabilities and establish a challenging experimental setup. Finally, we design a novel retrieval mechanism that leverages LLM feedback to select positive and negative candidates and employs a new training objective to prioritize historically significant sequences. This comprehensive framework bridges the gap between retrieval models and LLMs, enhancing the time-series forecasting process.
3.1. Dataset Construction
To address the limitations of existing datasets and better align with real-world economic analysis, we construct new datasets by integrating financial indicators alongside historical stock prices, enabling deeper insights into market behavior that stock prices alone cannot capture. This section details stock selection, raw data collection (Collection of Basic Price Indicators) and financial indicator selection process(Calculation and Selection of Financial Indicators). Specifically, our datasets are built using high-trade-volume U.S. stocks, while the indicators are selected through domain expert consultation and mutual information analysis.
Stock Selection. To construct our datasets, we select high-trade-volume U.S. stocks across three periods: 2014-2015, 2017-2018, and 2022-2023. The first two periods align with two benchmark datasets: for 2014-2015, we use the same stocks as the ACL18 dataset (Xu and Cohen, 2018), and for 2017-2018, we use the same stocks as the BIGDATA22 dataset (Soun et al., 2022). To incorporate more recent data, we manually select high-trade-volume stocks from 2022-2023 to create the STOCK23 dataset. This ensures our training and evaluation reflect recent market conditions and provide a robust benchmark for stock movement prediction.
Collection of Basic Price Indicators. Stock data is collected using the Yahoo Finance API (Xu and Berkely, 2014), which provides basic price indicators such as the opening price, highest price, lowest price, adjusted closing price, and trading volume for each trading day. The daily price movement is then calculated as the reference answer for model training and evaluation.
We then divide the stock data into training, validation, and test sets. To ensure temporal integrity and prevent information leakage, we partition the dataset by stocks rather than by time. This approach maintains consistent time ranges across splits while ensuring that stocks in the validation and test sets are entirely unseen during training. Unlike time-based splits, stock-based partitioning provides a more robust evaluation of the model’s ability to generalize to new stocks, better reflecting real-world scenarios where models are applied to unseen entities. Table 2 summarizes the stock counts for each split in our datasets.
Table 2. Dataset statistics.
Dataset Stock amount Trading dates
train valid test all sequences query sequences
ACL18 (Xu and Cohen, 2018) 33 5 33 2014.06.02-2015.12.31 2015.06.03-2015.12.31
BIGDATA22 (Soun et al., 2022) 22 3 22 2019.04.01-2020.12.31 2020.04.09-2020.12.31
STOCK23 24 3 24 2022.01.03-2023.12.31 2023.01.03-2023.12.31
Calculation and Selection of Financial Indicators. In our datasets, we integrate financial indicators to provide critical insights into market behavior that stock prices alone cannot capture. Following analyzing methods in the book Machine learning for trading (Jansen, 2020), we conduct an analysis using query sequences in all training sets to explore the relationship between commonly used financial indicators and returns. To measure the relationship between these indicators and returns, we utilize Mutual Information (MI), a non-linear measure of dependency that quantifies the amount of information one variable contains about another. Specifically, the MI between an indicator $X$ and returns $R_{t}$ is calculated as:
[…]
where $p(x,r)$ represents joint probability density function of $X$ (indicator) and $R_{t}$ (returns), $p(x)$ shows marginal probability density function of $X$, and $p(r)$ is marginal probability density function of $R_{t}$.
We compute MI scores for each indicator and select the top-20 indicator groups with the highest scores as our candidates, with the detailed score ranking provided in Appendix B.1. Additionally, an example of a data piece from our datastore is illustrated in Appendix A.1.
3.2. Sequence Serialization

### v3

# Retrieval-augmented Large Language Models for Financial Time Series Forecasting

@@
Abstract
Accurately forecasting stock price movements is critical for informed financial decision-making, supporting applications ranging from algorithmic trading to risk management. However, this task remains challenging due to the difficulty of retrieving subtle yet high-impact patterns from noisy financial time-series data, where conventional retrieval methods, whether based on generic language models or simplistic numeric similarity, often fail to capture the intricate temporal dependencies and context-specific signals essential for precise market prediction. To bridge this gap, we introduce FinSrag [^2] Code and data are available at FinSrag ., the first retrieval-augmented generation (RAG) framework with a novel domain-specific retriever FinSeer for financial time-series forecasting. FinSeer leverages a candidate selection mechanism refined by LLM feedback and a similarity-driven training objective to align queries with historically influential sequences while filtering out financial noise. Such training enables FinSeer to identify the most relevant time-series data segments for downstream forecasting tasks, unlike embedding or distance-based retrieval methods used in existing RAG frameworks. The retrieved patterns are then fed into StockLLM, a 1B-parameter LLM fine-tuned for stock movement prediction, which serves as the generative backbone. Beyond the retrieval method, we enrich the retrieval corpus by curating new datasets that integrate a broader set of financial indicators, capturing previously overlooked market dynamics. Experiments demonstrate that FinSeer outperforms existing textual retrievers and traditional distance-based retrieval approaches in enhancing the prediction accuracy of StockLLM, underscoring the importance of domain-specific retrieval frameworks in handling the complexity of financial time-series data.
1 Introduction
Financial time-series forecasting is crucial for maintaining market stability and efficiency, with profound implications for investment decisions, risk assessment, and economic policy formulation (Fama and French, 2000). However, the extreme volatility and nonlinear dynamics of financial markets pose significant analytical challenges, requiring sophisticated approaches to distill actionable signals from complex, noise-laden data streams. Stock movement prediction, which involves determining future price direction (up or down), represents a particularly critical yet demanding task that has attracted substantial research interest (Xie et al., 2023a, 2024a, 2024b). Although recent LLM-based approaches have shown promise in stock prediction, they predominantly rely on textual data like news and social media (Wang et al., 2024), using only the past several days’ closing prices as reference while largely ignoring the wealth of information contained in historical time-series patterns (Bustos and Pomares-Quimbaya, 2020). This oversight creates a critical gap in effectively harnessing temporal financial data for prediction. The challenge is further compounded by the enormous scale and complexity of such data, spanning multiple influencing factors and extended historical contexts. Addressing this requires an intelligent retrieval mechanism capable of efficiently navigating vast time-series datasets to extract and deliver the most relevant patterns, thereby empowering LLMs to generate more accurate and reliable stock movement forecasts.
Despite the critical importance of retrieving relevant historical patterns for stock movement prediction, current embedding-based and distance-based retrieval-augmented generation (RAG) approaches (Joshi et al., 2024; He et al., 2024; Cui et al., 2024) face fundamental limitations when applied to financial time-series data. Conventional embedding-based methods struggle because numeric financial sequences often exhibit superficially similar patterns that lack explicit semantic meaning, making it difficult for text-trained retrievers to distinguish meaningful signals from spurious correlations. Meanwhile, distance-based techniques like Dynamic Time Warping (DTW) (Yang et al., 2024) prove inadequate as they typically rely on single-variable comparisons (e.g., adjusted closing prices), ignoring the rich contextual information from other financial indicators. This approach limits retrieval to simplistic trend matching and fails to capture deeper, context-dependent relationships where sequences with divergent or opposing trends may actually contain complementary predictive signals. The inherent complexity of financial markets demands retrieval methods that can move beyond surface-level pattern matching to identify truly informative historical sequences that may exhibit complex, non-obvious relationships to current market conditions.
Figure 1: Overview of our time-series RAG framework.
To address these challenges, we introduce FinSrag, the first Financial time-Series RAG framework for stock movement prediction, using FinSeer, a Financial Time-Series Retriever to effectively retrieve the most beneficial historical sequences beyond surface-level pattern matching for downstream forecasting. FinSeer is trained to identify sequences that are embedded with essential or complementary information with limited similarity compared with the query that are beneficial for prediction. A key challenge in training FinSeer lies in the absence of explicit retrieval ground truth, that for any given query, it is inherently ambiguous which historical patterns will most effectively aid forecasting given the vast amount of data. To address this, we propose an LLM-guided relevance estimation, where the language model itself is used to assess and refine the selection of candidate sequences during training. Building on the approach of Zhang et al. (2023), we assess candidate relevance by feeding query-candidate pairs into the LLM and using the generation logits corresponding to the correct answer as a proxy for relevance. Intuitively, candidates yielding higher logits are considered more informative for the forecasting task.
We then train FinSeer to effectively distinguish high-value sequences (positives) from irrelevant sequences (negatives) by learning a retrieval embedding space where relevant candidates are closely aligned with their corresponding queries. Specifically, for each query, FinSeer is optimized to maximize the similarity score between the query embedding and the embedding of top-ranked candidates (i.e., those associated with higher generation logits from the LLM), while simultaneously pushing apart embeddings of low-ranked, uninformative candidates. This contrastive learning objective encourages FinSeer to internalize the LLM’s implicit relevance signals, thereby enhancing its ability to retrieve sequences that are truly beneficial for downstream forecasting.
Beyond improving relevant sequence retrieval, our framework introduces a crucial advancement in what to retrieve by expanding the candidate pool beyond traditional price-based segments. Specifically, we segment the financial time series into candidates, where each candidate corresponds to a time-series segment of a single feature, such as the adjusted close price or a specific technical indicator. Unlike prior approaches that retrieve only from price sequences, for the first time, our method incorporates 28 additional financial indicators, allowing retrieval over a richer set of candidate sequences that capture diverse aspects of market behavior [^3] For instance, candidates derived from the KDJ indicator group, particularly in the overbought region where $K>80$, $D>70$, and $J>90$ can reveal early signals of potential reversals (Wu and Diao, 2015).. To support this, we construct a new dataset and enriched existing two where time-series segments from both price data and a broad set of financial indicators are included as retrieval candidates, enabling FinSeer to discover deeper predictive patterns beyond surface-level price trends.
Comprehensive evaluations across multiple financial time-series datasets demonstrate that FinSeer, our specialized retriever, is the only approach that consistently enhances LLM forecasting accuracy, outperforming five state-of-the-art retrievers from the MTEB leaderboard when integrated into the FinSrag framework. This superior performance not only validates our domain-specific design for time-series data, but also crucially reveals FinSeer’s unique capability to identify and retrieve the most predictive financial indicator segments. Notably, the distance-based method which cannot leverage FinSrag’s architecture, shows weaker performance, further confirming the advantages of our framework’s structured approach to sequence selection, by incorporating diverse financial indicators as candidates.
More than just improving accuracy, FinSeer’s retrieval patterns provide valuable financial insights: By automatically discovering and leveraging latent market factors that conventional approaches miss, our method offers an empirically grounded alternative to theoretical assumptions in economic models. This data-driven approach enables systematic identification of predictive market features, opening new avenues to study market inefficiencies and dynamic dependencies through the lens of actual predictive relevance rather than ex-ante hypotheses.
In conclusion, our contributions are summarized as follows:
  1. 1.
    We introduce FinSrag, the first RAG framework with domain-specific retriever for financial time-series forecasting. By improving the capability of identifying and retrieving the most relevant time-series data segments, FinSrag significantly enhances LLMs’ forecasting capabilities in financial markets.
  2. 2.
    We propose FinSeer, a novel domain-specific retriever trained via LLM feedback to uncover overlooked market signals. We complement this with an enriched dataset featuring 28 expert-selected technical indicators that go beyond conventional price data.
  3. 3.
    Experimental results demonstrate the effectiveness of our framework, which surpasseses state-of-the-art retrievers and distance-based methods.
2 Prelimilaries
@@
Retrieval-augmented financial time-series forecasting (Sezer et al., 2020) involves predicting future values or trends ($G$) based on a given query sequence ($q$) and a set of retrieved historical sequences ($c$). These sequences are collected over time at regular intervals. In the retrieval process, the goal of the retrieval model ($R$) is to efficiently identify and extract the most useful historical sequences from a large pool of candidates. By providing relevant context, the retrieval model enhances the forecasting model’s ability to make accurate and reliable predictions. In the specific task of stock movement prediction (Xu and Cohen, 2018), the problem is framed as a binary classification task: predicting whether a stock’s price will rise or fall on the next trading day. Given a query sequence $q$, which represents the stock’s price over the previous $t$ days, the model retrieves relevant sequences as context and predicts the stock’s movement $M_{q,d}$ for the next trading day $d$. Details about threshold settings are illustrated in Appendix C.
3 Methodology
3.1 Retrieval Candidate Pool Design and Dataset Construction
To support our retrieval framework, we introduce a new dataset comprising high-volume stocks from 2022-2023 and enhance two existing datasets with 28 carefully selected financial indicators. Complete dataset statistics are presented in Table 5, while partitioning details and additional specifications appear in Appendix E.1. The following parts detail our methodology for: (1) stock price data collection and indicator selection (in Collection of Indicators), and (2) temporal scope determination and content specification for queries and candidate sequences (in Query and Candidate Construction).
Table 1: Dataset statistics of query and candidate sequence amounts [^5] In the table, the candidate rows represent the complete temporal range of each dataset. For any specific query, the corresponding candidate pool spans from the dataset’s start date through the trading day immediately preceding the query date, as the example shown in Figure 2(b)..
Dataset License Trading Dates Train Valid Test
ACL18 (Xu and Cohen, 2018) MIT License query 2015.06.03-2015.12.31 3,312 440 2,912
candidate 2014.06.02-2015.12.31 441,694 67,320 444,312
BIGDATA22 (Soun et al., 2022) Public query 2020.04.09-2020.12.31 3,229 434 2,868
candidate 2019.04.01-2020.12.31 328,372 44,778 328,372
STOCK23 MIT License query 2023.01.03-2023.12.31 4,268 570 4,128
candidate 2022.01.03-2023.12.31 404,736 50,592 404,736
Collection of Indicators. Our financial data foundation comes from the Yahoo Finance API (Xu and Berkely, 2014), which provides 6 daily trading metrics, including opening, highest, lowest, close prices, adjusted close prices, along with trading volume. While these basic price data offer essential market snapshots, they fail to capture deeper market dynamics crucial for forecasting. For example, if the adjusted close price increases on continuous trading dates, it only shows upward momentum but cannot determine when a trend is exhausted. The KDJ indicator, however, combines price momentum and range position to signal overbought (>80) or oversold (<20) conditions Wu and Diao (2015).
These limitations motivate our multi-layer dataset design to extract deeper and latent market signals beyond raw price movements. This multi-layer feature design is inspired by established practices in quantitative finance and machine learning  Heaton et al. (2017); Jansen (2020); Chen et al. (2024), which emphasize the importance of hierarchical signal extraction from raw market data to enhance model predictive power. The dataset consists of three hierarchical components: (1) Core Price Metrics: We begin with aforementioned 6 fundamental indicators, including raw OHLCV (Open, High, Low, Close, Volume) and Adjusted Close prices. (2) Primary Technical Indicators: To capture mid-level market dynamics (e.g., trend strength, volatility, and price-volume interactions) Lo et al. (2000); Park and Irwin (2007); Wu and Diao (2015), which are not directly observable in raw OHLCV, we incorporate 10 widely-used technical indicators such as log returns, price momentum, and VWAP, whose practice is supported by both empirical studies and industry applications Chan (2013); Fischer and Krauss (2018); Gu et al. (2020a); Ma and Yan (2022). (3) Alpha Factors: To uncover higher-order predictive signals, we identify 18 Alpha Factors via a Mutual Information (MI)-based selection process Guyon and Elisseeff (2003); Jansen (2020). This approach quantifies the nonlinear dependency between candidate features and future returns, allowing us to retain only those indicators with strong predictive relationships. The effectiveness of alpha factors in enhancing financial model performance has been extensively supported in the literature Harvey et al. (2016); Tulchinsky (2019); Gu et al. (2020b); Kakushadze (2016).
The integrated dataset, exemplified in Appendix E.3, combines these three layers (6 price metrics + 10 technical indicators + 18 alpha factors) to provide a multidimensional view of market conditions. Complete specifications for all 34 indicators appear in Table 4.
Query and Candidate Construction. The query construction process involves temporal range determination and query content specification. For temporal boundaries, we establish that each query must have sufficient candidate sequences available for retrieval by requiring that the query date occurs at least one year after the start date of the corresponding dataset split. This temporal buffer ensures adequate historical context for meaningful retrieval. We implement a one-day sliding window across trading days to generate consecutive queries. For the example shown in Figure 2(a), in the ACL18 dataset spanning from 2014-06-02 to 2015-12-31, we define queries as those occurring between 2015-06-03 and 2015-12-31. This configuration guarantees that even the earliest query (2015-06-03) can access a full year of preceding candidate sequences. Each query consists of the stock identifier, query date, and recent market data for query-candidate matching. The market data includes adjusted closing prices from the five most recent trading days. The five-day window aligns with standard financial practices, where multiples of five are commonly used for price change calculations.
The candidate construction follows similar temporal and content specifications as queries. Candidates are drawn from all available historical data preceding each query date, spanning from the dataset’s start date to the trading day before the query. We apply a one-day sliding window to generate sequential candidates, with all stocks in the dataset eligible as candidates regardless of query stock matching. Each candidate consists of the stock ticker, price movement direction on a specific trading date, a relevant financial indicator, and corresponding indicator values calculated from the preceding five consecutive trading days. Figure 2(b) demonstrates this construction through examples of both rising and falling candidates retrieved from the pool.
Our approach enables seamless market monitoring through incremental updates to the candidate pool. When advancing the query date (e.g., from 2015-12-30 to 2015-12-31), we simply append the new trading day’s data (2015-12-30) to the existing pool (2014-06-02 through 2015-12-29), maintaining all historical candidates without requiring model retraining. This efficient update mechanism ensures continuous operation while preserving the complete candidate history.
Figure 2: Illustration of query and candidate construction. (a) illustrates how to construct all queries for a given dataset. (b) illustrates the corresponding candidate pool for a given query, and the update of candidate pool with trading date updates.
3.1.1 Sequence Serialization